OpenStudy (anonymous):

How many years will it take for a $500 to grow to $1006.64 at 6% interest compounded quarterly? A) 11.50 years B) 11.75 years C) 12.00 years D) 12.25 years E) 12.50 years

OpenStudy (amistre64):

1006.64 = 500(1+.06/4)^(4*t)

OpenStudy (amistre64):

1006.64 ------- = [1+(.06/4)]^(4t) 500

OpenStudy (amistre64):

log(2.01328) = log(1.015)^4t log(2.01328) ----------- = 4t log(1.015) log(2.01328) ------------ = t 4 log(1.015)

OpenStudy (amistre64):

t = 11.75 years t = 11 yrs and 9 mths

OpenStudy (amistre64):

dble check by using this value of t in the original

OpenStudy (amistre64):

1006.64=500(1.015)^(4*11.75) 1006.54 = 500(1.015)^47 1006.54 = 500(2.01) 1006.54 = 1005 ...because i chopped the decimal off to much lol

OpenStudy (amistre64):

1006.6395513857092257218559330282 when you dont truncate it lol